February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Contents

Foreword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v

1 Proton Structure

1.1 B. Portheault: High Q^{2} Structure Functions and Parton

Distributions . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 D. Kcira: Structure Functions and the Transition Region

from Photoproduction to Deep Inelastic Scattering . . . . . 15 1.3 S. Moch, J.A.M. Vermaseren and A. Vogt: Deep Inelastic

Structure Functions at NNLO and Beyond . . . . . . . . . 26 1.4 V. Chekelian: Gluon Density and Strong Coupling from

Structure Functions and Measurement of F_{L} . . . . . . . . 34 1.5 S. Alekhin: Power corrections in the Deep Inelastic Scatter-

ing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 1.6 J. Pumplin: The Art of Global Fitting for Parton Distribu-

tions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

2 Spin Physics 2.1 M. Tytgat: New Results from HERMES . . . . . . . . . . 67 2.2 D. Boer: Theoretical Aspects of Spin Physics . . . . . . . 77 2.3 U. Stösslein: Effects of e^{±} Polarization on Final States at

HERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

3 Production of Hadrons, Jets and Photons

3.1 J. Cvach: Photoproduction of Jets and Prompt Photons at

HERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 3.2 G. Heinrich: Photoproduction of Isolated Photons, Single

Hadrons and Jets at NLO . . . . . . . . . . . . . . . . . . . 114 3.3 C. Glasman: Jet Production in Deep Inelastic ep Scattering

at HERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 3.4 L. Jönsson: Forward Jet and Particle Production at Low x 136 3.5 M. Maniatis: Single Hadron Production in Deep Inelastic

Scattering . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 3.6 F. Corriveau: Photo- and Electroproduction of Single

Hadrons and Resonances . . . . . . . . . . . . . . . . . . . 158 3.7 P. Minkowski and W. Ochs: Gluonic Meson Production . 169

4 Heavy-Flavour Production 4.1 S. Padhi: Open Charm and Beauty Production . . . . . . 183 4.2 I. Schienbein: Open Heavy-Flavour Photoproduction at

NLO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 4.3 P. Thompson: Experimental Results on Exclusive Vector

Mesons and Heavy Quarkonium . . . . . . . . . . . . . . . 207

5 Elastic and Diffractive ep Scattering 5.1 S. Levonian: Diffractive Structure Functions and QCD Fits 219 5.2 N. Vlasov: Diffractive Jet and Charm Production . . . . . 231 5.3 R. B. Peschanski: Beyond BFKL . . . . . . . . . . . . . . 241 5.4 O. Nachtmann: Pomeron Physics and QCD . . . . . . . . 253 5.5 L. Motyka: Theory of Diffractive Meson Production … 268 5.6 G. Gustafson: Unintegrated Parton Densities and Applica-

tions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 5.7 F. Schrempp and A. Utermann: Instanton-Driven Satura-

tion at Small x . . . . . . . . . . . . . . . . . . . . . . . . . 289

6 New Physics at HERA

6.1 K. O. Diener: QCD Corrections to Single Top Quark and

W Boson Production at HERA . . . . . . . . . . . . . . . . 305 6.2 E. Gallo: Search for Beyond the Standard Model Physics

at HERA . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315

7 Plans for HERA-II and Beyond

7.1 E.C. Aschenauer: Plans for the HERMES Experiment at

HERA II . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327 7.2 M. Klein: Physics with H1 at HERA II . . . . . . . . . . 330 7.3 R. Yoshida: Prospects and Status of ZEUS at HERA II . 347 7.4 A. C. Caldwell: A New Round of Experiments for HERA 352

List of participants 365 Foreword

The international workshop entitled New Trends in HERA Physics 2003 took place from 28 September to 3 October 2003 at Ringberg Castle, which overlooks Lake Tegernsee in the foothills of the Bavarian Alps, one of the most picturesque locations to be found in the whole of Germany. The castle was built during the first half of the twentieth century by Duke Luitpold in Bavaria (Herzog Luitpold in Bayern), a member of the Wittelsbach family which ruled Bavaria over 800 years, and his friend Friedrich Attenhuber, an all-round artist, architect, and interior decorator. The castle is entirely their creation, from the massive Renaissance-inspired exterior right down to the fittings and furniture, which, in every detail, were designed by Attenhuber himself and executed by native craftsmen. Attenhuber also painted every single picture exhibited in the castle. He found his models in the farmhouses around Lake Tegernsee. The castle embodies all trends of art and styles which dominated the first half of last century, combined with local Alpine originality and the individual creative power of its constructors. According to the Duke’s last will, the castle passed into the hands of the Max Planck Society after his death, in 1973. The castle was then transformed into a conference venue, where scientists can exchange their latest ideas and discuss problems with their colleagues from all over the world in beautiful surroundings and in a relaxed mountain atmosphere, high above the daily business activities.

This was the fourth event in a series of Ringberg workshops on HERA physics, which was started in 1997 and was continued in 1999 and 2001. In fact, at the end of these workshops, many participants expressed the opinion that this was a successful endeavour to bring theorists and experimentalists together in order to interpret the latest HERA data, and that it would be useful to organize a follow-up workshop in the same spirit.

On the occasion of the 2003 Ringberg workshop, forty-one experts of elementary-particle physics, both theorists and experimentalists, from twenty-three universities and research institutions in fourteen countries congregated to present their latest results on the various aspects of HERA physics. Specifically, there were sixteen theorists and twenty-five experimentalists, the latter representing the H1, HERMES, and ZEUS collaborations at HERA. The topics included: proton structure function; polarized ep scattering; photoproduction of hadrons, jets, and prompt photons; final states in deep inelastic scattering, with special emphasis on hadrons, jets, resonances, and gluonic mesons; heavy-flavour and charmonium production; elastic and diffractive ep scattering; new physics at HERA; and plans for HERA-II and beyond. We hope that the high-energy-physics community will benefit from these proceedings, in which the ongoing efforts in understanding the nature of the strong interactions, with particular emphasis on HERA physics, are documented.

We wish to thank all our friends and colleagues who have contributed to these proceedings. We are indebted to the workshop secretary, Mrs. Rosita Jurgeleit, for her assistance before, during, and after the workshop and to Dr. Annette Holtkamp for her technical assistance in the editorial work. The local costs at Ringberg Castle and the costs for the publication and dissemination of these proceedings were covered in equal parts by the Deutsches Elektronen-Synchrotron at Hamburg and the Max-Planck-Institut für Physik at Munich, which we gratefully acknowledge.

Hamburg, January 2004 Günter Grindhammer Bernd A. Kniehl Gustav Kramer Wolfgang Ochs

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 1

Proton Structure

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

1.1. B. Portheault: High Q^{2} Structure Functions and Parton Distributions 1.2. D. Kcira: Structure Functions and the Transition

Region from Photoproduction to Deep Inelastic Scattering

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

1.3. S. Moch, J.A.M. Vermaseren and A. Vogt: Deep Inelastic Structure Functions at NNLO and Beyond

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

1.4. V. Chekelian: Gluon Density and Strong Coupling from Structure Functions and Measurement of F_{L}

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

1.5. S. Alekhin: Power corrections in the Deep Inelastic Scattering

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

1.6. J. Pumplin: The Art of Global Fitting for Parton Distributions

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 2

Spin Physics

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

2.1. M. Tytgat: New Results from HERMES

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

2.2. D. Boer: Theoretical Aspects of Spin Physics

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

2.3. U. Stösslein: Effects of e^{±} Polarization on Final States at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 3

Production of Hadrons, Jets and Photons

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.1. J. Cvach: Photoproduction of Jets and Prompt Photons at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.2. G. Heinrich: Photoproduction of Isolated Photons, Single Hadrons and Jets at NLO

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.3. C. Glasman: Jet Production in Deep Inelastic ep Scattering at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.4. L. Jönsson: Forward Jet and Particle Production at Low x

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.5. M. Maniatis: Single Hadron Production in Deep Inelastic Scattering

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.6. F. Corriveau: Photo- and Electroproduction of Single Hadrons and Resonances

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

3.7. P. Minkowski and W. Ochs: Gluonic Meson Production

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 4

Heavy-Flavour Production

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

4.1. S. Padhi: Open Charm and Beauty Production

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

4.2. I. Schienbein: Open Heavy-Flavour Photoproduction at NLO

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

4.3. P. Thompson: Experimental Results on Exclusive Vector Mesons and Heavy Quarkonium

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 5

Elastic and Diffractive ep Scattering

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.1. S. Levonian: Diffractive Structure Functions and QCD Fits

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.2. N. Vlasov: Diffractive Jet and Charm Production

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.3. R. B. Peschanski: Beyond BFKL

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.4. O. Nachtmann: Pomeron Physics and QCD

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.5. L. Motyka: Theory of Diffractive Meson Production

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.6. G. Gustafson: Unintegrated Parton Densities and Applications

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

5.7. F. Schrempp and A. Utermann: Instanton-Driven Saturation at Small x

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 6

New Physics at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

6.1. K. O. Diener: QCD Corrections to Single Top Quark and W Boson Production at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

6.2. E. Gallo: Search for Beyond the Standard Model Physics at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

Chapter 7

Plans for HERA-II and Beyond

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

7.1. E.C. Aschenauer: Plans for the HERMES Experiment at HERA II

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

7.2. M. Klein: Physics with H1 at HERA II

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

7.3. R. Yoshida: Prospects and Status of ZEUS at HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

7.4. A. C. Caldwell: A New Round of Experiments for HERA

February 13, 2004 12:37 WSPC/Trim Size: 9in x 6in for Proceedings 03con

List of participants

Sergey Alekhin  IHEP Protvino  alekhin@sirius.ihep.su
Elke-Caroline Aschenauer DESY Hamburg  elke.aschenauer@desy.de
Daniel Boer  U Amsterdam  dboer@nat.vu.nl

Yuraj Bracinik  MPI Munich  juraj.bracinik@desy.de

Gerd Buschhorn  MPI Munich  gwb@mppmu.mpg.de
Allen Caldwell  MPI Munich  caldwell@mppmu.mpg.de
Vladimir Chekelian  MPI Munich  shekeln@mail.desy.de
Francois Corriveau  U McGill  francois@mail.desy.de
Jaroslav Cvach  Acad. of Sci. Prague cvach@fzu.cz
Kai Diener  PSI Villigen  kai.diener@psi.ch
Elisabetta Gallo  U Florence, INFN  gallo@mail.desy.de
Claudia Glasman  U Madrid  claudia@mail.desy.de
Günter Grindhammer  MPI Munich  guenterg@desy.de
Gösta Gustafson  U Lund  gosta@thep.lu.se

Gudrun Heinrich  Durham, IPPP  gudrun.heinrich@durham.ac.uk
Leif Jönsson  U Lund  leif.jonsson@hep.lu.se

Rosita Jurgeleit  MPI Munich  roj@mppmu.mpg.de
Dorian Kcira  U Wisconsin  kcira@mail.desy.de
Christian Kiesling  MPI Munich  cmk@mppmu.mpg.de
Max Klein  DESY Zeuthen  klein@ifh.de

Bernd Kniehl  U Hamburg  kniehl@mail.desy.de

Gustav Kramer  U Hamburg  kramer@mail.desy.de
Sergey Levonian  DESY Hamburg  levonian@mail.desy.de
Markos Maniatis  U Hamburg  markos@mail.desy.de
Sven Moch  DESY Zeuthen  moch@ifh.de
Leszek Motyka  Jagellonian U Cracow leszekm@th.if.uj.edu.pl

Otto Nachtmann  U Heidelberg  o.nachtmann@thphys.uni-heidelberg.de

Wolfgang Ochs  MPI Munich  wwo@mppmu.mpg.de
Sanjay Padhi  DESY Hamburg  spadhi@mail.desy.de
Robert Peschanski  CNRS Saclay  pesch@spht.saclay.cea.fr
Ringaile Placakyte  MPI Munich  ringaile@mppmu.mpg.de
Benjamin Portheault  LAL Orsay  portheau@lal.in2p3.fr
Jon Pumplin  Michigan State U  pumplin@pa.msu.edu
Zuzana Rurikova  MPI Hamburg  zuzana.rurikova@desy.de
Ingo Schienbein  DESY Hamburg  schien@mail.desy.de
Fridger Schrempp  DESY Hamburg  fridger.schrempp@desy.de
Uta Stösslein  DESY Hamburg  uta.stoesslein@desy.de
Paul Thompson  U Birmingham  thompspd@mail.desy.de

Michael Tytgat  U Gent  michael@inwfsun1.UGent.be

Nikolai Vlasov  U Freiburg  vlasov@mail.desy.de
Biljana Vujicic  MPI Munich  bivuj@mppmu.mpg.de

Rik Yoshida  Argonne  rik.yoshida@desy.de